OpenStudy (anonymous):

how do you do this 1 :( sec(arctan2)

OpenStudy (anonymous):

i have to get exact value for this

OpenStudy (anonymous):

it's the sqrt of 5

OpenStudy (anonymous):

how did you get the answer :D

OpenStudy (dumbcow):

sec^2 = tan^2 +1 sec(x) = sqrt(tan^2 +1) sqrt(tan(arctan2)^2 +1) = sqrt(5)

OpenStudy (anonymous):

ty